For the following reaction<br /> ClO₄⁻(aq) + 2H⁺(aq) + Pb²⁺(aq) ⇄ ClO₃⁻(aq) + H₂O(l) + Pb⁴⁺(aq) K=1 × 10⁻¹⁵<br /> which of the following is true?

Question 3 options:

ΔG° > 0 the reactants are more stable than the products.

ΔG° > 0 the products are more stable than the reactants.

ΔG° < 0 the reactants are more stable than the products.

ΔG° < 0 the products are more stable than the reactants.

ΔG° = 0 the reactants and products are equally stable
Answer #1

clou (aq) + 2H+ (aq) + Pb²+ (aq) & co (aq) + H₂O(1) + Phut (aq) -15 K - X10 R = 8.314 T/K mol 25°c = 25+273 = 298 K T AG = -